I need help solving the following equation.
solve for x.
4x(x-1)-5x(x)=3

Answer by sachi(392)

I need help solving the following equation.
solve for x.
4x(x-1)-5x(x)=3
ans:-
4x^2-4x-5x^2=3
or x^2+4x+3=0
or x^2+x+3x+3=0
or x(x+1)+3(x+1)=0
or (x+1)(x+3)=0
or x=-1 or -3 ans